OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). 11 19 16 8 i will fan and medal

OpenStudy (alexandervonhumboldt2):

substitute 4 for x and solve

TheSmartOne (thesmartone):

plug in x = 4

OpenStudy (alexandervonhumboldt2):

g(4)=4^2+3 solve

TheSmartOne (thesmartone):

Is your question \(\sf g(x) = x^2 + 3\) or what?

OpenStudy (alexandervonhumboldt2):

yes it is

TheSmartOne (thesmartone):

you never know, so don't make assumptions :)

OpenStudy (anonymous):

SO is it c

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

4*4=16 16+3=?

OpenStudy (anonymous):

b??

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

thank you!!!